OpenStudy (anonymous):

Basic problem necessary for future understanding of Schrodinger's Equation: A wave has the form y=Acos(2πx/λ+π/3)when x<0. For x > 0, the wavelength is λ/2. By applying continuity conditions at x = 0, find the amplitude (in terms of A) and phase of the wave in the region x > 0. Sketch the wave, showing both x < 0 and x>0.

OpenStudy (anonymous):

This is much more a mathematical problem: Create a condition such that: \[ y(x)= \begin{cases} A\cos\left(\frac{2\pi x}{\lambda}+\frac{\pi}{3}\right)\text{ when } x<0\\ A\cos\left(\frac{\pi x}{\lambda}+\frac{\pi}{3}\right)\text{ when } x>0 \end{cases} \]Is continuous at \(x=0\)

OpenStudy (anonymous):

And find the amplitude, etc, etc. Although, I'm not 100% sure of the phase, but I believe the above to be the case.

OpenStudy (anonymous):

Be mindful of arithmetic ^_^ \[y(x)= \left\{ \begin{array}{l} A\cos\Big(\frac{2 \pi x}{\lambda} + \frac{\pi}{3}\Big) \text{ when} \ x<0 \\ A \cos \Big( \frac{4 \pi x}{\lambda} + \delta\Big) \text{ when} \ x>0 \end{array} \right.\]